# custom_references.bib

### v1

@@ 5.1 Protocols with Longer Messages
For $k\geq 2$, we proceed by using the same exact algorithm as in Theorem 5.1 - for $\alpha\in\{0,1\}^{r_{1}}\times\dots\times\{0,1\}^{r_{k-1}}$, we view $\pi_{\alpha}:(\{0,1\}^{r_{k}})^{\ell}$ as a function over $\ell r_{k}$ bits. We set some parameters $c,h$ and find sets $B_{R},B_{H},B_{I}$ where $B_{R},B_{H}$ will find bad bits for the last round and $B_{I}$ will apply induction to find bad players that can corrupt the first $i-1$ rounds. When applying induction for $(k-1)$-round protocols, we call it with parameter $\delta_{I}$ that we set later. So, our inductive calls to $k-1$ round protocol will be such that it corrupts at most $\delta_{I}\ell$ bad players. We will then choose bad players for the last round such that they will control the bits that are in $B_{R}\cup B_{H}$. We see that in that case, the number of bad players we control over last round is at most $\left\lvert B_{R}\right\rvert+\left\lvert B_{H}\right\rvert$.
@@
Finally, we bound $\left\lvert B_{I}\right\rvert$. To do that, we first observe that by choice of $h$, we have that $h>\frac{\log(h)}{10^{6}}$ and $h>\frac{\log(1/\delta)}{10^{6}}$. So, $42\log(4h/\delta)<h/$ 4. Using this, we see:
@@
where in the last inequality we used the bound $42\log(4h/\delta)<h/$ 4. Lastly, we show that we can indeed set $\delta_{I}$ to the desired value, i.e., that $\varepsilon\geq\frac{C\log(1/\gamma\delta_{I})\log^{(k+1)}(\ell)}{\log^{(k)}(\ell)}$. To show this inequality holds, we will make use of the fact that $\varepsilon\geq\frac{C\log(1/\gamma\delta)\log^{(k+2)}(\ell)}{\log^{(k+1)}(\ell)}$. We see that
@@
We apply induction on $k\geq 1$. For $k=1$, the result follows by [KKL88influence]. For $k\geq 2$, we proceed as follows. We will maintain $2$ kinds of bad players: First, for $1\leq i\leq k$, let $B^{(i)}_{R},B^{(i)}_{H}\subset[\ell]$ be the set of bad players that we will use to corrupt round $i$. Second, for $1\leq i\leq k$, we let $B_{I}^{(i)}\subset[\ell]$ be the set of bad players that we will use to corrupt players for rounds $1$ to $i$ (these will be provided to us by induction). At the end, we will take the union of all these sets, and that will be the final set of bad players that we will use to corrupt $\pi$. Let $S_{k}$ be the set of messages across $k$ rounds that we will eventually be able to corrupt. Initially, $S_{k}=\left(\{0,1\}^{\ell}\right)^{k}$ and as the algorithm proceeds we will restrict it so tht at each step, the set of messages in $S_{k}$ are the current candidate set of messages that we could possibly corrupt. For $11\leq i\leq k-$ 1, let $S_{i}$ be the projection of $S_{k}$ on the first $i$ coordinates. Often we will remove a string $s$ from $S_{i}$. That means we will remove all strings from $S_{k}$ whose projection on the first $i$ coordinates is $s$, and update $S_{1},\dots,S_{k-1}$ so that they accurately hold the projections of $S_{k}$. Also, for $1\leq i\leq k$, let $\sigma_{i}=\frac{\left\lvert S_{i}\right\rvert}{2^{i\ell-\sum_{j=1}^{i}\left(\left\lvert B^{(j)}_{R}\right\rvert+\left\lvert B^{(j)}_{H}\right\rvert+\left\lvert B_{I}^{(j)}\right\rvert\right)}}$, the density of strings in $S_{i}$. Let $h^{(1)},\dots,h^{(k)},c^{(1)},\dots,c^{(k)}\in[\ell]$ be parameters that we set later.
@@
      For each $\alpha\in S_{k-1}$, let $\pi_{\alpha}^{(k)}:\{0,1\}^{\ell}\to\{0,1\}$ be the induced round $k$ protocol. We keep string $\alpha\in S_{k-1}$ iff $\Pr[\pi_{\alpha}^{(k)}=1]\geq\gamma/k$. Using reverse Markov argument from 3.3, since $\Pr[\pi=1]\geq\gamma$, we have that after this operation, $\sigma_{k-1}\geq\gamma/4$.
@@
      We now try to simultaneously corrupt them last round functions. Set =$\mathcal{F}=\{\pi_{\alpha}^{(k)}\}_{\alpha\in S_{k-1}}$ and apply Theorem 4.1 with parameters $h=h^{(k)},\delta=1/3$. Let $B_{R}^{(k)}$ be the random set, and for $\alpha\in\left(\{0,1\}^{\ell}\right)^{k-2}$ let $g_{\alpha}^{(k-1)}:\{0,1\}^{\ell}\to\left([\ell]\cup 0\right)^{h^{(k)}}\cup\bot$ encode the heavy set (of size $h^{(k)}$) so that $B_{R}^{(k)}\cup g_{\alpha}^{(k-1)}$ can $(1-\gamma/k)$-bias $\pi_{\alpha}^{(k)}$ towards $1$ (if possible for that $\alpha$; otherwise $g$ outputs $\bot$).
@@
      Let $\pi_{temp}$ be $k-1$ round protocol such that $\pi_{temp}(x)=1$ iff $x\in S_{k-1}$. We see that $\Pr[\pi_{temp}=1]\geq\gamma/6$. By induction, we can find set of bad players, say $B_{I}^{(k-1)}$, so that $\Pr[\pi_{temp}|_{B_{I}^{(k-1)}}=1]\geq 2/3$. Restrict $S_{k-1}$ to only these assignments so that $\sigma_{k-1}\geq 2/3$.
@@
    For $i$ going down from $k-1$ to $1$, proceed as follows:
@@
      Our goal will be to simultaneously corrupt the maps $g_{\alpha}^{(i)}:\{0,1\}^{\ell}\to([\ell]\cup 0)^{h^{(i+1)}}\cup\bot$ for $\alpha\in S_{i-1}$. Let $C_{1},\dots,C_{h^{(i+1)}}\subset[\ell]$ be such that for all $\alpha\in S_{i}$, for all $j\in[h^{(i+1)}]$ the output of $g_{\alpha,j}^{(i)}$ lies in $C_{j}$ with high probability. Initially, $C_{1},\dots,C_{h^{(i+1)}}$ are set to equal $[\ell]$. We now corrupt players in round $i$ so that each of the $C_{j}$ will have $\left\lvert C_{j}\right\rvert\leq c^{(i)}$.
@@ 6 Constructing Improved Constant-Round Coin Flipping Protocols
To help construct our protocol, we need to introduce the following notion of an assembly:
@@
    (Grouping) For any $t\geq 1$, in 00 rounds, $A$ can be transformed into a ${\left(b,n/t,st\right)}$-assembly.
@@
    (Splitting) For any $t\geq 1$, in 00 rounds, $A$ can be transformed into a ${\left(bt,nt,s/t\right)}$-assembly.
@@
    (Feige’s Lightest Bin Protocol [feige99lightestbin]) For any $\beta\leq 2^{s},0<\delta<1$, in $1$ round, $A$ can be transformed into a ${\left(\frac{b}{\beta}+\frac{\delta(n-b)}{\beta},\frac{n}{\beta},s\right)}$-assembly, with success probability at least $1-\beta\cdot\exp\left(\frac{-\delta^{2}(n-b)}{2\beta}\right)$
@@
Lemma 6.4 (Resilient function from [AL93resilient, IV25resilient]).
@@
[…]
@@
    For rounds $2$ to $k-1$, we proceed as follows: In round $i$, we are given a ${\left(\frac{\gamma(\log^{(i-1)}(\ell))^{2}}{(\log^{(k)}(\ell))^{2}}+(\log^{(i-1)}(\ell))^{1.76},\frac{(\log^{(i-1)}(\ell))^{2}}{3},3\log^{(i-1)}(\ell)\right)}$-assembly.
@@
By a union bound over all $k$ rounds (the failure probability is dominated by the $i=k-1$ iteration of step 2(b) above), we get that all our transformations succeed with probability $1-\exp(-(\log^{(k-1)}(\ell))^{1.48})$. In such an event, the error from our final round - round $k$ from applying the resilient function is $O\left(\gamma+(\log^{(k-1)}(\ell))^{-0.23}\right)$. Since $(\log^{(k-1)}(\ell))^{-0.23}>\exp(-(\log^{(k-1)}(\ell))^{1.48})$, we get that the final error of our coin flip is at most $O\left(\gamma+(\log^{(k-1)}(\ell))^{-0.2}\right)$ as desired. ∎
@@
    To do this, we split each set $S_{i}$ arbitrarily into $t$ parts. We indeed obtain $tn$ sets of size $s/t$ each. Moreover, whenever we split a set consisting of only good players, all of the resulting sets only consist of good players. Hence, the number of such sets in the resultant assembly is $(n-b)t=nt-bt$ as desired.
@@
    To do this, we use Feige’s lightest bin protocol from [feige99lightestbin]. In particular, we set up $\beta$ bins $B_{1},\dots,B_{\beta}$ which receive ‘votes’ as follows. In one round, the players in each set $S_{1},\dots,S_{n}$ each flip a random coin. For each set $S_{i}$, we interpret their $s$ coin flips together as generating a random number between $1$ and $\beta$ (since $\beta\leq 2^{s}$) and that number is the ‘vote’ for which bin the set should go to. After this voting process, we pick the lightest bin (the one that has received the fewest votes), say $B_{i^{*}}$, and let the sets in that bin (along with arbitrary other sets, chosen in some canonical way so that number of sets in the output assembly is $n/\beta$) form the output assembly .
    We analyze the process described above. We treat all players in any bad set as bad, i.e., they are allowed to cast their votes after seeing the votes of all the good sets. We first note that the number of sets in the lightest bin $B_{i^{*}}$ is at most $n/\beta$ (since there are $n$ sets and $\beta$ bins). Hence, the above process always outputs an assembly consisting of $n/\beta$ many sets of size $s$ each. We now lower bound the number of good players in each of the bins. Fix any bin $i\in[\beta]$ and let $\mathbf{X}_{i}$ be the random variable representing the number of good sets that voted $i$. We think of the vote of each of the good sets as independently choosing whether to vote for bin $i$ or not, where the vote is cast for bin $i$ with probability $1/\beta$. Hence $\operatorname*{\mathbb{E}}[\mathbf{X}_{i}]=(n-b)/\beta$, and by the Chernoff bound 3.4, we have that

### v2

@@ 5.1 Protocols with Longer Messages
For $k\geq 2$, we proceed by using the same exact algorithm as in Theorem 5.1 - for $\alpha\in\{0,1\}^{r_{1}}\times\dots\times\{0,1\}^{r_{k-1}}$, we view $\pi_{\alpha}:(\{0,1\}^{r_{k}})^{\ell}$ as a function over $\ell r_{k}$ bits. We set some parameters $c,h$ and find sets $B_{R},B_{H},B_{I}$ where $B_{R},B_{H}$ will find bad bits for the last round and $B_{I}$ will apply induction to find bad players that can corrupt the first i1i-1 rounds. When applying induction for (k1)(k-1)-round protocols, we call it with parameter $\delta_{I}$ that we set later. So, our inductive calls to k1k-1 round protocol will be such that it corrupts at most $\delta_{I}\ell$ bad players. We will then choose bad players for the last round such that they will control the bits that are in $B_{R}\cup B_{H}$. We see that in that case, the number of bad players we control over last round is at most $\left\lvert B_{R}\right\rvert+\left\lvert B_{H}\right\rvert$.
@@
Finally, we bound $\left\lvert B_{I}\right\rvert$. To do that, we first observe that by choice of $h$, we have that $h>\frac{\log(h)}{10^{6}}$ and $h>\frac{\log(1/\delta)}{10^{6}}$. So, $2\log(4h/\delta)<h/4$. Using this, we see:
@@
where in the last inequality we used the bound $2\log(4h/\delta)<h/4$. Lastly, we show that we can indeed set $\delta_{I}$ to the desired value, i.e., that $\varepsilon\geq\frac{C\log(1/\gamma\delta_{I})\log^{(k+1)}(\ell)}{\log^{(k)}(\ell)}$. To show this inequality holds, we will make use of the fact that $\varepsilon\geq\frac{C\log(1/\gamma\delta)\log^{(k+2)}(\ell)}{\log^{(k+1)}(\ell)}$. We see that
@@
5.2 Biasing leader election protocols
We here show how to bias $k$-round leader election protocols as specified in Corollary 5.3. We do this by repeated applications of Theorem 5.1.
Proof of Corollary 5.3.
We fix a leader election protocol $\pi$. We will iteratively decrease the output domain of $\pi$ by a factor of half, each time invoking Corollary 5.3 while doing so. Let $H_{0}=[\ell]$ and $B_{0}=\varnothing$. We proceed in $t=\log^{(k+1)}(\ell)$ steps. At the end of each step $i$ we will maintain the invariant that $\Pr[\pi|_{B_{i}}\in H_{i}]\geq 1-\gamma$. For $i=1,\dots,t$, we proceed as follows:
  1. 1.
    The invariant property guarantees us that $\Pr[\pi|_{B_{i-1}}\in H_{i-1}]\geq 1-\gamma$. So by an averaging argument, there exists $H_{i}\subset H_{i-1}$ with $\left\lvert H_{i}\right\rvert=\left\lvert H_{i-1}\right\rvert/2$ such that $\Pr[\pi|_{B_{i-1}}\in H_{i-1}]\geq(1-\gamma)/2\geq\gamma$.
  2. 2.
    We let $\pi^{(i)}$ be the $k$-round coin flipping protocol over $\ell-\left\lvert B_{i-1}\right\rvert$ players that on the message transcript $\alpha$, outputs $1$ if $\pi|_{B_{i-1}}(\alpha)\in H_{i}$ and outputs 0 otherwise (here $C_{0}$ is a universal constant).
  3. 3.
    Since $\Pr[\pi|_{B_{i-1}}\in H_{i-1}]\geq\gamma$, we have that $\Pr[\pi^{(i)}=1]\geq\gamma$. We apply Theorem 5.1 to find bad players $B\subset[\ell]\setminus B_{i-1}$ with $\left\lvert B\right\rvert\leq\frac{C_{0}(\ell-\left\lvert B_{i-1}\right\rvert)}{\gamma\log^{(k)}(\ell-\left\lvert B_{i-1}\right\rvert)}$ such that it can bias $\pi^{(i)}$ to output $1$ with probability $1-\gamma$.
  4. 4.
    We let $B_{i}=B\cup B_{i-1}$ where $B$ is from the previous step.
  5. 5.
    We then indeed see that $\Pr[\pi|_{B_{i}}\in H_{i}]\geq 1-\gamma$, maintaining the invariant property.
We let the final set of bad players be $B_{t}\cup H_{t}$. By the invariant property at the end of step $t$, we do have that $\pi$ outputs from $H_{t}$ when corrupted by $B_{t}$ with probability $1-\gamma$. Since now both are part of the set of bad players, we conclude that a bad player is indeed elected the leader with probability $\geq 1-\gamma$.
It remains to bound the number of bad players. We individually bound $H_{t}$ and $B_{t}$ as follows: Since $\left\lvert H_{i}\right\rvert=\left\lvert H_{i-1}\right\rvert/2$ for all steps $i$, we easily see that $\left\lvert H_{t}\right\rvert=\left\lvert H_{0}\right\rvert/2^{t}=\ell/2^{t}=\frac{\ell}{\log^{(k)}(\ell)}$. To bound $\left\lvert B_{t}\right\rvert$, we first observe that the size of the set of bad players $B$ added in each step is upper bounded by $\frac{C_{0}(\ell-\left\lvert B_{i-1}\right\rvert)}{\gamma\log^{(k)}(\ell-\left\lvert B_{i-1}\right\rvert)}\leq\frac{C_{0}(\ell)}{\gamma\log^{(k)}(\ell)}$ Using this, we easily conclude that $\left\lvert B_{t}\right\rvert\leq t\cdot\frac{C_{0}(\ell)}{\gamma\log^{(k)}(\ell)}=\log^{(k+1)}(\ell)\cdot\frac{C_{0}(\ell)}{\gamma\log^{(k)}(\ell)}$.
Thus, the total number of bad players is at most $\log^{(k+1)}(\ell)\cdot\frac{2C_{0}(\ell)}{\gamma\log^{(k)}(\ell)}$. Setting our universal constant to be $2C_{0}$, we obtain the desired claim. ∎
Remark 5.9.
We could also have obtained $1$ round leader election lower bounds directly from Theorem 8.1 by setting $\Delta=\log(\log(\ell))$. Our parameter dependence would have been slightly worse - we would have needed $\frac{\ell\log(\log(\ell))^{2}\operatorname{poly}(\log(\log(\log(\ell))))}{\log(\ell)}$ many bad players had we done so. This would still be a doubly exponential improvement over all previous $1$-round leader election lower bounds. Moreover, our random selection lower bounds from Theorem 8.1 only rely on the Edge Isoperimetric inequality instead of the lower bound on influence due to KKL [KKL88] as above.
@@ 5.2 Biasing leader election protocols
We apply induction on $k\geq 1$. For $k=1$, the result follows by [KKL88]. For $k\geq 2$, we proceed as follows. We will maintain $2$ kinds of bad players: First, for $1\leq i\leq k$, let $B^{(i)}_{R},B^{(i)}_{H}\subset$$\ell$] be the set of bad players that we will use to corrupt round $i$. Second, for $1\leq i\leq k$, we let $B_{I}^{(i)}\subset[\ell]$ be the set of bad players that we will use to corrupt players for rounds $1$ to $i$ (these will be provided to us by induction). At the end, we will take the union of all these sets, and that will be the final set of bad players that we will use to corrupt $\pi$. Let $S_{k}$ be the set of messages across $k$ rounds that we will eventually be able to corrupt. Initially, $S_{k}=\left(\{0,1\}^{\ell}\right)^{k}$ and as the algorithm proceeds we will restrict it so tht at each step, the set of messages in $S_{k}$ are the current candidate set of messages that we could possibly corrupt. For $1\leq i\leq k-1$, let $S_{i}$ be the projection of $S_{k}$ on the first $i$ coordinates. Often we will remove a string $s$ from $S_{i}$. That means we will remove all strings from $S_{k}$ whose projection on the first $i$ coordinates is $s$, and update $S_{1},\dots,S_{k-1}$ so that they accurately hold the projections of $S_{k}$. Also, for $1\leq i\leq k$, let $\sigma_{i}=\frac{\left\lvert S_{i}\right\rvert}{2^{i\ell-\sum_{j=1}^{i}\left(\left\lvert B^{(j)}_{R}\right\rvert+\left\lvert B^{(j)}_{H}\right\rvert+\left\lvert B_{I}^{(j)}\right\rvert\right)}}$, the density of strings in $S_{i}$. Let $h^{(1)},\dots,h^{(k)},c^{(1)},\dots,c^{(k)}\in[\ell]$ be parameters that we set later.
@@
      For each $\alpha\in S_{k-1}$, let $\pi_{\alpha}^{(k)}:\{0,1\}^{\ell}\to\{0,1\}$ be the induced round $k$ protocol. We keep string $\alpha\in S_{k-1}$ iff $\Pr[\pi_{\alpha}^{(k)}=1]\geq\gamma/k$. Using reverse Markov argument from 3.4, since $\Pr[\pi=1]\geq\gamma$, we have that after this operation, $\sigma_{k-1}\geq\gamma/4$.
@@
      We now try to simultaneously corrupt them last round functions. Set $\mathcal{F}=\{\pi_{\alpha}^{(k)}\}_{\alpha\in S_{k-1}}$ and apply Theorem 4.1 with parameters $h=h^{(k)},\delta=1/3$. Let $B_{R}^{(k)}$ be the random set, and for $\alpha\in\left(\{0,1\}^{\ell}\right)^{k-2}$ let $g_{\alpha}^{(k-1)}:\{0,1\}^{\ell}\to\left([\ell]\cup 0\right)^{h^{(k)}}\cup\bot$ encode the heavy set (of size $h^{(k)}$) so that $B_{R}^{(k)}\cup g_{\alpha}^{(k-1)}$ can $(1-\gamma/k)$-bias $\pi_{\alpha}^{(k)}$ towards $1$ (if possible for that $\alpha$; otherwise $g$ outputs $\bot$).
@@
      Let $\pi_{temp}$ be k1k-1 round protocol such that $\pi_{temp}(x)=1$ iff $x\in S_{k-1}$. We see that $\Pr[\pi_{temp}=1]\geq\gamma/6$. By induction, we can find set of bad players, say $B_{I}^{(k-1)}$, so that $\Pr[\pi_{temp}|_{B_{I}^{(k-1)}}=1]\geq 2/3$. Restrict $S_{k-1}$ to only these assignments so that $\sigma_{k-1}\geq 2/3$.
@@
    For $i$ going down from k1k-1 to $1$, proceed as follows:
@@
      Our goal will be to simultaneously corrupt the maps $g_{\alpha}^{(i)}:\{0,1\}^{\ell}\to([\ell]\cup 0)^{h^{(i+1)}}\cup\bot$ for $\alpha\in S_{i-1}$. Let $C_{1},\dots,C_{h^{(i+1)}}\subset[\ell]$ be such that for all $\alpha\in S_{i}$, for all $j\in[h^{(i+1)}]$ the output of $g_{\alpha,j}^{(i)}$ lies in $C_{j}$ with high probability. Initially, $C_{1},\dots,C_{h^{(i+1)}}$ are set to equal $\ell$]. We now corrupt players in round $i$ so that each of the $C_{j}$ will have $\left\lvert C_{j}\right\rvert\leq c^{(i)}$.
@@ 6 Constructing Improved Constant-Round Coin Flipping Protocols
To help construct our protocol, we need to introduce the notion of an assembly:
@@
    (Grouping) For any $t\geq 1$, in 0 rounds, $A$ can be transformed into a $\left(b,n/t,st\right)$}-assembly.
@@
    (Splitting) For any $t\geq 1$, in 0 rounds, $A$ can be transformed into a $\left(bt,nt,s/t\right)$}-assembly.
@@
    (Feige’s Lightest Bin Protocol [Fei99]) For any $\beta\leq 2^{s},0<\delta<1$, in $1$ round, $A$ can be transformed into a $\left(\frac{b}{\beta}+\frac{\delta(n-b)}{\beta},\frac{n}{\beta},s\right)$}-assembly, with success probability at least $1-\beta\cdot\exp\left(\frac{-\delta^{2}(n-b)}{2\beta}\right)$
@@
Lemma 6.4 (Resilient function from [AL93, IV24]).
@@
      We use transformation 3 from Lemma 6.3 (Feige’s lightest bin protocol) with $\beta=\frac{\ell}{(\log(\ell))^{3}},\delta=(\log(\ell))^{-0.25}$ to transform the ${\left(\frac{\gamma\ell}{\log(\ell)\log^{(k)}(\ell)},\frac{\ell}{3\log(\ell)},3\log(\ell)\right)}\textrm{-assembly}$ into a $\left(\frac{\gamma(\log(\ell))^{2}}{(\log^{(k)}(\ell))^{2}}+(\log(\ell))^{1.76},\frac{(\log(\ell))^{2}}{3},3\log(\ell)\right)$}-assembly, succeeding with probability $\geq 1-\ell\exp\left(-(\log(\ell))^{1.49}\right)$.
@@
    For rounds $2$ to k1k-1, we proceed as follows: In round $i$, we are given a $\left(\frac{\gamma(\log^{(i-1)}(\ell))^{2}}{(\log^{(k)}(\ell))^{2}}+(\log^{(i-1)}(\ell))^{1.76},\frac{(\log^{(i-1)}(\ell))^{2}}{3},3\log^{(i-1)}(\ell)\right)$}-assembly.
@@
By a union bound over all $k$ rounds (the failure probability is dominated by the i=k1i=k-1 iteration of step 2(b) above), we get that all our transformations succeed with probability $1-\exp(-(\log^{(k-1)}(\ell))^{1.48})$. In such an event, the error from our final round - round $k$ from applying the resilient function is $O\left(\gamma+(\log^{(k-1)}(\ell))^{-0.23}\right)$. Since $(\log^{(k-1)}(\ell))^{-0.23}>\exp(-(\log^{(k-1)}(\ell))^{1.48})$, we get that the final error of our coin flip is at most $O\left(\gamma+(\log^{(k-1)}(\ell))^{-0.2}\right)$ as desired. ∎
@@
    To do this, we split each set $S_{i}$ arbitrarily into $t$ parts. We indeed obtain $tn$ sets of size $s/t$ each. Moreover, whenever we split a set consisting of only good players, all of the resulting sets only consist of good players. Hence, the number of such sets in the resultant assembly is (nb)t=ntbt(n-b)t=nt-bt as desired.
@@
    To do this, we use Feige’s lightest bin protocol from [Fei99]. In particular, we set up $\beta$ bins $B_{1},\dots,B_{\beta}$ which receive ‘votes’ as follows. In one round, the players in each set $S_{1},\dots,S_{n}$ each flip a random coin. For each set $S_{i}$, we interpret their $s$ coin flips together as generating a random number between $1$ and $\beta$ (since $\beta\leq 2^{s}$) and that number is the ‘vote’ for which bin the set should go to. After this voting process, we pick the lightest bin (the one that has received the fewest votes), say $B_{i^{*}}$, and let the sets in that bin (along with arbitrary other sets, chosen in some canonical way so that number of sets in the output assembly is $n/\beta$) form the output assembly .
    We analyze the process described above. We treat all players in any bad set as bad, i.e., they are allowed to cast their votes after seeing the votes of all the good sets. We first note that the number of sets in the lightest bin $B_{i^{*}}$ is at most $n/\beta$ (since there are $n$ sets and $\beta$ bins). Hence, the above process always outputs an assembly consisting of $n/\beta$ many sets of size $s$ each. We now lower bound the number of good players in each of the bins. Fix any bin $i\in[\beta]$ and let $\mathbf{X}_{i}$ be the random variable representing the number of good sets that voted $i$. We think of the vote of each of the good sets as independently choosing whether to vote for bin $i$ or not, where the vote is cast for bin $i$ with probability $1/\beta$. Hence $\operatorname*{\mathbb{E}}[\mathbf{X}_{i}]=(n-b)/\beta$, and by the Chernoff bound 3.5, we have that